I need to find the perimeter of a right triangle that has a height of 9 and hypotenuse of 15. Please advise. Thank you.
Found 2 solutions by richwmiller, Edwin McCravy:
Answer by richwmiller(17219)

Not hypertense but hypotenuse
A triangle has three sides. We know two of them.
How do we find the third side of a right triangle?
a^2+b^2+c^2
where c is the hypotenuse and a and b are the legs.
a+b+c=perimeter
so first we need to find the missing side.
a^2+9^2=15^2
a^2+81=225
a^2=144
a=12
We could also have deduced the third side since this is a multiple of the classic triplet right triangle 3, 4,5
3*3=9
3*4=12
3*5=15
So now you have the three sides.
You still need to add them together to get the perimeter
a+b+c=perimeter

Answer by Edwin McCravy(20056)



We need to find x by the Pythagorean theorem:

(bottom leg)2 + (right leg)2 = (hypotenuse)2

                     x2 + 92 = 152

                     x2 + 81 = 225

                          x2 = 225 - 81

                          x2 = 144

                           x = 12

So the perimeter is 15 + 9 + 12 = 36  

Edwin